Giving a test to a group of students, the grades and gender are summarized below A B C Total Male 11 16 18 45 Female 13 20 10 43 Total 24 36 28 88 Let π represent the percentage of all male students who would receive a grade of A on this test. Use a 99.5% confidence interval to estimate p to three decimal places. Enter your answer as a tri-linear inequality using decimals (not percents).
Answer #1

I have answered the question below

Please up vote for the same and thanks!!!

Do reach out in the comments for any queries

Answer:

Prob that male students getting A grade = 11/45 = 0.2444

Calculate or Op = vp^(1 - pa) VN

= sqrt(0.244(1- 0.244)/45) = 0.06406

Z score for 99.5% = 2.807

Calculate high end confidence interval total: High End = p^+ zscorea X Op

= 0.2444 + 2.807* 0.06406 = 0.424

Calculate low end confidence interval total: Low End = p^ - Zscoreç x 0p

= 0.2444 - 2.807* 0.06406 = 0.065

=> 0.065<p< 0.424